# Noise-Conditioned Mixture-of-Experts Framework for Robust Speaker Verification

### v2

@@ Abstract
Robust speaker verification under noisy conditions remains an open challenge. Conventional deep learning methods learn a robust unified speaker representation space against diverse background noise and achieve significant improvement. In contrast, this paper presents a noise-conditioned mixture-of-experts framework that decomposes the feature space into specialized noise-aware subspaces for speaker verification. Specifically, we propose a noise-conditioned expert routing mechanism, a universal model based expert specialization strategy, and an SNR-decaying curriculum learning protocol, collectively improving model robustness and generalization under diverse noise conditions. The proposed method can automatically route inputs to expert networks based on noise information derived from the inputs, where each expert targets distinct noise characteristics while preserving speaker identity information. Comprehensive experiments demonstrate consistent superiority over baselines, confirming that explicit noise-dependent feature modeling significantly enhances robustness without sacrificing verification accuracy.
@@ I Introduction
To alleviate the above-mentioned problem, a dominant and effective approach is to employ speech enhancement (SE) networks for noise suppression, with the enhanced feature then utilized for SV [12, 3]. Compared to cascading pre-trained SE and SV networks, [14] showed that constructing an SE model specialized for SV tasks yields better performance. Thereafter, some studies have attempted to boost system robustness through targeted improvements to SE modules. In [22], both masking- and mapping-based SE networks are integrated into the SV system to remove noise from different aspects. In [8], a novel extened U-Net is adopted as the SE backbone and showed the superiority of the jointly optimized cascade system through end-to-end learning. With the advent of diffusion models, recent works [9][10] have incorporated them into the SE front-end, achieving notable performance gains.
@@
While both kinds of approaches have shown promising results, their dependence on unified feature-space modeling may present certain constraints. In cases where input distributions exhibit notable variations, maintaining effective discrimination within a single feature space could prove challenging. Building upon this, we explore a noise-conditioned mixture-of-expert (NCMoE) framework and investigate the feasibility of decomposing the feature space into noise-specific subspaces. This exploration into condition-specific subspace modeling is intended to offer a potential alternative for speaker feature extraction in challenging environments, and we hope it may provide useful insights for robust speaker representation learning. Specifically, a lightweight convolutional network first estimates the noise characteristics of spectral features to guide the selection of specialized expert branches. Subsequently, to facilitate expert training, we introduce a universal model-based strategy. This strategy begins with a generalist model and then progressively specializes it into noise-adapted subspaces, a process further stabilized by an SNR-progressive curriculum. Experiments on VoxCeleb1 with simulated noise conditions suggest this approach outperforms conventional baselines.
@@ II-A Overview of the Framework
As illustrated in Fig.1 (a), our framework preserves the backbone network’s original architecture while augmenting a selected intermediate layer with parallel expert branches to balance model capacity and efficiency. Each expert replicates the complete structure of the original layer (i.e., maintaining identical input/output dimensions and internal connections), ensuring architectural consistency. In addition, a compact noise classification network (Fig.1 (b)) dynamically selects a single expert branch per sample during forward propagation, keeping others inactive. This design ensures computational efficiency through sparse expert activation while maintaining manageable resource requirements, allowing specialized processing tailored to distinct noise conditions. In practice, this work instantiates the strategy by augmenting the second residual stage of a ResNet into multiple experts. The core design is generic and can be transplanted to other backbone networks.
@@ III-B Implementation Details
The acoustic front-end extracts 80-dimensional log-mel filterbank features. The experiments are based on the ResNet34 backbone with 32 initial channels, and the embedding dimension is 256. The model is trained for 150 epochs using SGD optimizer with the AAM-Softmax loss for speaker classification, employing mixed precision training to accelerate computation while maintaining stability. The proposed framework extends this baseline with four parallel residual branches in residual stage two, each following the identical structure as mentioned in Section II-2.1, while the noise classifier maintains a 32-channel initial width and adheres to the protocol of doubling channels during time-frequency downsampling. We treat reverberation as a structured noise and accordingly define four distinct augmentation categories as noise-type labels: “noise”, “babble”, “music” and “reverberation”. The temperature factor $\gamma$ is empirically set at 0.1 , with the truncated Gaussian distribution bounded between 20 dB and 0 dB, and $\sigma$ empirically set at 0.2. The coefficient $k$ in Eq.3 is derived as 7.6 by logarithmic transformation to ensure that $\mu$ decays smoothly from 20 to 0.01 (approximately 0 for numerical stability) during training. The implementation is based on the Wespeaker [18] toolkit using 2 NVIDIA RTX 5070 Ti GPUs.
@@ III-D Further Analysis
TABLE III: ABLATION STUDY RESULTS (AVERAGE EER% ACROSS 5 SNR LEVELS) UNDER VARIOUS SYNTHETIC NOISE CONDITIONS.
@@
Our ablation study in Table III demonstrates varying degrees of performance degradation when removing three key components: the UMES, NCER, and SDCL. The most significant performance drop occurs with the removal of the UMES strategy, indicating that excessive specialization among experts compromises discriminative speaker feature extraction due to their inherent shared patterns. Although eliminating the noise classification loss (w/o NCER) degrades performance, the model still maintains superiority over the baseline. Furthermore, the ablation of SDCL confirms that progressively reducing SNR during training effectively enhances model robustness.
@@
Figure 3: Confusion Matrix of EER Under Diverse Training and Testing Audio Conditions.
To evaluate the performance of individual experts in our mixture-of-experts framework under varied test conditions, each expert (based on ECAPA-TDNN) was tested on data at 0 dB SNR. The resulting confusion matrix is shown in Fig.3. For clearer visualization of performance shifts under train-test mismatch, the experts were independently initialized without weight sharing. Results show that performance degrades under most mismatched conditions, with the babble expert exhibiting the most severe degradation when tested on nonspeech interference. In contrast, both noise and music experts maintain relatively stable performance under mismatch and also perform well on nonspeech interference.

### v3

@@ Abstract
Robust speaker verification under noisy conditions remains an open challenge. Conventional deep learning methods learn a robust unified speaker representation space against diverse background noise and achieve significant improvement. In contrast, this paper presents a noise-conditioned mixture-of-experts framework that decomposes the feature space into specialized noise-aware subspaces for speaker verification. Specifically, we propose a noise-conditioned expert routing mechanism, a universal model based expert specialization strategy, and an SNR-decaying curriculum learning protocol, collectively improving model robustness and generalization under diverse noise conditions. The proposed method can automatically route inputs to expert networks based on noise information derived from the inputs, where each expert targets distinct noise characteristics while preserving speaker identity information. Comprehensive experiments demonstrate consistent superiority over baselines.
@@ I Introduction
To alleviate the above-mentioned problem, a dominant and effective approach is to employ speech enhancement (SE) networks for noise suppression, with the enhanced feature then utilized for SV [16, 4]. Compared to cascading pre-trained SE and SV networks, [19] showed that constructing an SE model specialized for SV tasks yields better performance. Thereafter, some studies have attempted to boost system robustness through targeted improvements to SE modules. In [27], both masking- and mapping-based SE networks are integrated into the SV system to remove noise from different aspects. In [11], a novel extened U-Net is adopted as the SE backbone and showed the superiority of the jointly optimized cascade system through end-to-end learning. To mitigate error accumulation, [13] aggregates lost information for speaker restoration, while [6] preserves speaker information via explicit noise modeling with dual U-Nets. With the advent of diffusion models, recent works [12][14] have incorporated them into the SE front-end, achieving notable performance gains.
@@
While both kinds of approaches have shown promising results, their dependence on unified feature-space modeling may present certain constraints. In cases where input distributions exhibit notable variations, maintaining effective discrimination within a single feature space could prove challenging. Mixture-of-experts (MoE) framework [9, 18], which leverage multiple specialized sub-networks, have demonstrated strong performance across various tasks by enabling more flexible feature representations [2]. However, their potential for improving robustness in speaker verification remains underexplored. Building upon this, we explore a noise-conditioned mixture-of-expert (NCMoE) framework and investigate the feasibility of decomposing the feature space into noise-specific subspaces. This exploration into condition-specific subspace modeling is intended to offer a potential alternative for speaker feature extraction in challenging environments, and we hope it may provide useful insights for robust speaker representation learning. Specifically, a lightweight convolutional network first estimates the noise characteristics of spectral features to guide the selection of specialized expert branches. Subsequently, to facilitate expert training, we introduce a universal model-based strategy. This strategy begins with a generalist model and then progressively specializes it into noise-adapted subspaces, a process further stabilized by an SNR-progressive curriculum. Experiments on VoxCeleb1 with simulated noise conditions suggest this approach outperforms conventional baselines.
@@ II-A Overview of the Framework
As illustrated in Fig.1 (a), our framework preserves the backbone network’s original architecture while augmenting a selected intermediate layer with parallel expert branches to balance model capacity and efficiency. Each expert replicates the complete structure of the original layer (i.e., maintaining identical input/output dimensions and internal connections), ensuring architectural consistency. In addition, a compact noise classification network (Fig.1 (b)) dynamically selects a single expert branch per sample during forward propagation, keeping others inactive. This design ensures computational efficiency through sparse expert activation while maintaining manageable resource requirements, allowing specialized processing tailored to distinct noise conditions.
@@ III-B Implementation Details
The acoustic front-end extracts 80-dimensional log-mel filterbank features. Given its wide adoption, strong performance, and scalability, we use a ResNet with 32 initial channels as the baseline to validate the NCMoE framework. The speaker embedding dimension is 256. The model is trained for 150 epochs using SGD optimizer with the AAM-Softmax loss for speaker classification, employing mixed precision training to accelerate computation while maintaining stability. The proposed framework extends this baseline with four parallel residual branches in residual stage two, each following the identical structure as mentioned in Section II-2.1, while the noise classifier maintains a 32-channel initial width and adheres to the protocol of doubling channels during time-frequency downsampling. We treat reverberation as a structured noise and accordingly define four distinct augmentation categories as noise-type labels: “noise”, “babble”, “music” and “reverberation”. The temperature factor $\gamma$ is empirically set at 0.1 , with the truncated Gaussian distribution bounded between 20 dB and 0 dB, and $\sigma$ empirically set at 0.2. The coefficient $k$ in Eq.3 is derived as 7.6 by logarithmic transformation to ensure that $\mu$ decays smoothly from 20 to 0.01 (approximately 0 for numerical stability) during training. The implementation is based on the Wespeaker [23] toolkit using 2 NVIDIA RTX 5070 Ti GPUs.
@@ III-D Further Analysis
TABLE III: Ablation study results (average eer% across 5 snr levels) under varous synthetic noise conditions.
@@
TABLE IV: Sensitivity analysis (average eer% across 5 snr levels) on hyper-parameters ($\gamma$ and $\sigma$).
Setting $\gamma$=0.1 $\gamma$=1.0 $\sigma$=0.1 $\sigma$=0.2 $\sigma$=0.3 $\sigma$=1.0
Average 3.41 3.45 3.44 3.41 3.42 3.50
Our ablation study in Table III demonstrates varying degrees of performance degradation when removing three key components: the UMES, NCER, and SDCL. The most significant performance drop occurs with the removal of the UMES strategy, indicating that excessive specialization among experts compromises discriminative speaker feature extraction due to their inherent shared patterns. Although eliminating the noise classification loss (w/o NCER) degrades performance, the model still maintains superiority over the baseline. Furthermore, the ablation of SDCL confirms that progressively reducing SNR during training effectively enhances model robustness. Moreover, Table IV summarizes the impact of key hyper-parameters, showing that the results do not fluctuate significantly with their variations.
@@
Figure 3: Confusion Matrix of the routing result of the noise classifier and the EER of NCMoE under diverse training and testing audio conditions.
To analysis the behavior of the noise classifier and the performance of individual experts in NCMoE, each expert (based on ECAPA-TDNN) was tested on data at 0 dB SNR. The resulting confusion matrix is shown in Fig.3. Note that experts were independently initialized here for a clearer visualization. It shows that the noise classifier maintains high routing accuracy for babble and music, with moderate accuracy for noise. Out-of-domain data show distinct routing patterns, confirming the router’s discriminative capability and validating the NCMoE framework. In addition, results show that performance degrades under most mismatched conditions. We attribute this to the distinct spectral characteristics of speech, music, and noise, each necessitating a specific noise suppression pattern. Mismatched expert routing applies the wrong suppression mode, leading to performance degradation.